# Quantum diffeomorphisms cannot make indefinite causal order definite

### v2

@@ Abstract
The field of indefinite causal order has seen significant advancements in recent years. While classically the causal order of two timelike separated events A and B is fixed - either A before B or B before A - this is no longer true in quantum theory. There, it is possible to encounter superpositions of causal orders. The quantum switch is one of the most prominent processes with indefinite causal order. While the optical quantum switch has been successfully implemented in experiments, some argue that this merely simulates a process with indefinite causal order and that a superposition of spacetime metrics is required for a true realization. Here, we provide a relativistic definition of causal order, show that it encompasses both the optical and gravitational quantum switch, and does not differentiate between them. Moreover, we show that this notion of causal order is invariant under so-called quantum diffeomorphisms and that it is an operationally meaningful observable in both the general relativistic and quantum mechanical sense. Importantly, this observable does not distinguish between the indefinite causal order implemented in the optical and gravitational quantum switch, thus supporting the thesis that the optical quantum switch is just as much a realization of indefinite causal order as its gravitational counterpart.
The study of indefinite causal order (ICO) has seen remarkable progress over the last years both theoretically [1, 2, 3, 4] and experimentally [5, 6, 7, 8, 9, 10]. While indefinite causal order can arise in quantum theory alone, the field was inspired by the goal to combine general relativity and quantum theory [1]. General relativity is a deterministic theory with a dynamical causal structure. Quantum theory is a probabilistic theory with a fixed causal structure. Thus, combining the two would allow for processes to which no definite causal ordering can be assigned.
One of the simplest and most prominent implementations of ICO is the quantum switch [2, 3]. Depending on the state of a control qubit, two operations $U_{A}$ and $U_{B}$ are applied to a target system in a superposition of ‘$U_{A}$ before $U_{B}$’ and ‘$U_{B}$ before $U_{A}$’. Concretely, denoting the initial state of the target system by $\mathinner{|{\psi}\rangle}_{T}$ and taking the operations to be unitary, the resulting state is
@@
Various ways of implementing ICO have been reported and described in the literature. In the optical quantum switch, the target system is passing through the setup in a superposition of paths [5, 6, 7]. This type of ICO has been successfully implemented on the optical bench [8, 9, 10]. A different implementation is provided by the gravitational quantum switch [4, 12]. In this gedankenexperiment, a massive object in superposition of two locations sources a gravitational field in superposition, while two agents perform operations on a test particle. A minimal assumption of most quantum gravity theories is that the spacetime metric sourced by a mass in superposition is itself in superposition. Then, choosing an appropriate setup can result in an indefinite causal order due to the different time dilations experienced in each branch of the superposition.
@@
The radically different physics involved in the optical and the gravitational implementations of the quantum switch leads some to question whether it can be claimed that ‘genuine’ or ‘proper’ ICO has been experimentally realized. The quantum information community is roughly divided along two theses [5, 13, 12, 14, 15, 16]. On one hand, those that claim that the optical switch “can be seen as the first realization of a superposition of causal orders” [5], and on the other, those who view it as a mere simulation of ICO. The latter camp argues that “the current quantum switch experimental implementations do not feature superpositions of causal orders between spacetime events”, and that “these superpositions can only occur in the context of superposed gravitational fields” [12].
@@
To bring clarity into this debate, we cast the notion of ICO in a general relativistic language. This allows us to study the indefiniteness of causal order in a unified setting for both variations of the quantum switch. We first make precise the definition of an event in the context of a superposition of spacetimes. We then identify a single binary quantity that encodes the causal order of a process, the sign of the proper time, and demonstrate that it is invariant under quantum controlled diffeomorphisms — a quantum controlled arbitrary and independent change of coordinates in each branch of the superposition. We further show that this quantity is not only an operationally meaningful observable in the general relativistic but also in the quantum mechanical sense. Crucially, this observable does not differentiate whether ICO is arising due to a superposition of paths on a definite metric or a superposition of spacetimes, that is, it cannot distinguish between the optical and the gravitational quantum switch. This strengthens the claim that the optical and gravitational implementation of the quantum switch exhibit the same ‘type’ of ICO.
@@
This work is motivated by recent research on quantum reference frames [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], which suggests that the invariance of physical laws under coordinate transformations extends to quantum superpositions thereof. Indefinite causal order as we define it here is a prototype case of a physically meaningful quantity that has no classical counterpart and is an invariant under quantum coordinate changes.
@@
Take two systems and a test particle interacting with each of the systems once. Each interaction defines an event. As an example, consider two laboratories and an electron travelling between them. The events may correspond to the application of a unitary operation acting on the electron spin. We model this as three timelike curves $\gamma_{0}$, $\gamma_{A}$, and $\gamma_{B}$, with $\gamma_{0}$ coinciding once with $\gamma_{A}$ and once with $\gamma_{B}$ at the coincidences $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$ (see Fig. 1). Coincidences of worldlines define events in a coordinate independent way.
@@
$\tau_{\gamma}=\frac{1}{c}\int_{p_{\gamma,i}}^{p_{\gamma,f}}\sqrt{-g_{\mu\nu}dx^{\mu}dx^{\nu}},$ (2)
elapsed between two points $p_{\gamma,i}$ and $p_{\gamma,f}$. We take $p_{\gamma_{0},i}$ as the event at which the test particle is released and $p_{\gamma_{0},f}$ as the one at which a final measurement is applied. This fixes a time orientation along $\gamma_{0}$. ${\mathcal{E}_{B}}$ lies either in the future or past lightcone of ${\mathcal{E}_{A}}$. With our time orientation, ${\mathcal{E}_{B}}$ lies in the future of ${\mathcal{E}_{A}}$ if $\tau_{A}<\tau_{B}$. This corresponds to being able to signal from ${\mathcal{E}_{A}}$ to ${\mathcal{E}_{B}}$, which is also the operational notion of causal order employed in quantum information. Therefore, the causal order is encoded in the sign of the proper time difference. Define the crossing points $p_{\mathcal{E}_{A}}\equiv\gamma_{0}(\tau_{A})$, $p_{\mathcal{E}_{B}}\equiv\gamma_{0}(\tau_{B})$. The distinction between the events $\mathcal{E}_{X}$ and points $p_{\mathcal{E}_{X}}$ will be important in what follows.
@@
Consider two events $\mathcal{E}_{X}$, $X\in\{A,B\}$, connected by a timelike worldline $\gamma_{0}$ with crossing points $p_{\mathcal{E}_{X}}\equiv\gamma_{0}(\tau_{X})$. The causal order between ${\mathcal{E}_{A}}$ and ${\mathcal{E}_{B}}$ is the sign $s$ of the proper time difference $\Delta\tau=s|\tau_{B}-\tau_{A}|$.
Events on a superposition of spacetimes — Consider two globally hyperbolic spacetimes $g^{(1)}$ and $g^{(2)}$ each associated to a manifold $\mathcal{M}_{i}$, $i\in\{1,2\}$, diffeomorphic as differentiable manifolds. We take the same setup as above on each spacetime, with two events $\mathcal{E}_{A}^{(i)}$, $\mathcal{E}_{B}^{(i)}$ defined as coincidences. Next, we consider a superposition of two semi-classical, nearly orthogonal states of the metric, each peaked around $g^{(i)}$. Formally, using a control qubit whose orthogonal states $\mathinner{|{1}\rangle}$ and $\mathinner{|{2}\rangle}$ distinguish the branches, the quantum state is
@@
with $|\alpha|^{2}+|\beta|^{2}=1$  [^1] Note that Eq. (3) should be understood as suggestive since the Hilbert space assigned to the metric has not properly been introduced. However, if classically distinguishable spacetimes are represented by orthogonal states, one can form linear superpositions thereof and define a suitable inner product in the Hilbert space. They would correspond to the superposition of semi-classical states in a complete theory.. Such a scenario might, for example, arise when placing a gravitating object in a superposition of two locations [32, 33, 34, 35, 36, 4, 29].
In order to characterize the causal order between two events on a superposition of spacetimes, we first need to define the notion of an event itself in this context. Since $\mathcal{E}_{X}^{(1)}$ and $\mathcal{E}_{X}^{(2)}$, $X=A,B$ both represent the crossing of the worldline of the test particle with that of system $X$, we posit that they represent the same physical event:
@@
This extends the relativistic notion of coincidences to incorporate superpositions of trajectories.
Figure 2: A superposition of two scenarios with different causal orders $s^{(1)}$ and $s^{(2)}$. In each branch, we consider a manifold $\mathcal{M}_{1,2}$, a metric $g^{(1),(2)}$, and the worldlines of three systems. The black, blue, and red line depict $\gamma_{0}$, $\gamma_{A}$, and $\gamma_{B}$, with different transparencies illustrating the superposition. There are two events $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$ defined by the crossing of the test particle with one of the other trajectories. First, we apply a quantum-controlled diffeomorphism $\phi^{(1)}:\mathcal{M}_{1}\to\mathcal{M}$ together with a diffeomorphism $\phi^{(2)}:\mathcal{M}_{1}\to\mathcal{M}$ chosen such that each of the events is associated with a single point on $\mathcal{M}$. The metric and causal order may still be in superposition.
Note that it differs essentially from the notion of an event as a point on a spacetime manifold. This becomes particularly clear when counting the events involved in our scenario. There are four different manifold-points involved in our scenario – $p_{\mathcal{E}_{A}}^{(1)}$ and $p_{\mathcal{E}_{B}}^{(1)}$ on manifold $\mathcal{M}_{1}$ as well as $p_{\mathcal{E}_{A}}^{(2)}$ and $p_{\mathcal{E}_{B}}^{(2)}$ on manifold $\mathcal{M}_{2}$ (see Fig. 2). However, there are only two distinct physical events $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$ – the crossing of the test particle with system $A$ and the crossing of the test particle with system $B$. Causal order on a superposition of spacetimes — We can now extend Definition 1 of causal order between events to superpositions of spacetimes.
@@
Thus, the notions of definite and indefinite causal order are conveniently encoded in the product $s^{(1)}s^{(2)}$. Importantly, ICO in the above sense can be due to a superposition of spacetime metrics $g_{\mu\nu}(x)$ or of crossing points $p_{{\mathcal{E}_{A}}}$ and $p_{{\mathcal{E}_{B}}}$. In the following, we will refer to the latter as ICO due to delocalized events as, in this case, ${\mathcal{E}_{A}}$ and ${\mathcal{E}_{B}}$ are assigned different manifold points across the different branches. The above observation follows straightforwardly from Eq. (2), which tells us that the proper time difference depends on both the metric along the curve $\gamma_{0}$ and the coincidences of worldlines that define the events ${\mathcal{E}_{A}}$ and ${\mathcal{E}_{B}}$. Thus, this notion of ICO captures both the optical and the gravitational quantum switch.
@@
This does not mean, of course, that there are no other observables that can distinguish between the two scenarios. Were one to measure the curvature scalar, for example, one would find that it is zero in the optical and non-zero in the gravitational quantum switch in the realization of [4]. Nevertheless, having clarified what we mean by ICO in a general relativistic language, we find that both implementations lead to the same indefiniteness of causal order. In this sense, the optical quantum switch can be seen as just as much a realization of ICO as a hypothetical realization of the gravitational quantum switch.
@@
Note, however, that whether a process displays ICO ‘due’ to delocalized events or a superposition of spacetime metrics can in general change under quantum diffeomorphisms. To see this, consider a scenario with ICO. Let us assume that we begin in a quantum coordinate system in which ICO is due to delocalized events while the spacetime background is fixed. More specifically, we assume that the crossing points $p_{{\mathcal{E}_{A}}}^{(1)},p_{{\mathcal{E}_{B}}}^{(1)}$ and $p_{{\mathcal{E}_{A}}}^{(2)},p_{{\mathcal{E}_{B}}}^{(2)}$ differ across the branches of the superposition. Now, we can always find a quantum diffeomorphism that maps both configurations onto the same manifold $\mathcal{M}$ while ensuring that the events ${\mathcal{E}_{A}}$ and ${\mathcal{E}_{B}}$ are each associated with a single point on this manifold. That is, there always exist diffeomorphisms $\phi^{(1)}:\mathcal{M}_{1}\to\mathcal{M}$ and $\phi^{(2)}:\mathcal{M}_{2}\to\mathcal{M}$ such that
@@
Thus, whether the events ${\mathcal{E}_{A}}$ and ${\mathcal{E}_{B}}$ are localized or delocalized depends on the choice of quantum coordinates. If we had defined events in terms of spacetime points (as suggested e.g. in [12]), their number would change under a quantum diffeomorphism. But, as we have seen, the causal order itself is invariant under quantum diffeomorphisms; therefore, the setup must still exhibit ICO. This is because a quantum diffeomorphism also affects the metric. If we start with a definite spacetime metric $g$ and apply a different diffeomorphism in each branch, we end up with a superposition of spacetime metrics $g^{(i)}=(\phi^{(i)})^{-1}_{\ast}(g)$, $i\in\{1,2\}$ [^3] Note that the action of the diffeomorphisms on the fields is the inverse of their actions on the manifold points.. In this sense, a quantum diffeomorphism can shift the indefiniteness due to delocalized events into an indefiniteness of the spacetime metric at localized events. A concrete example of this transformation is provided in Appendix A.1.
Conversely, if we start with a superposition of different spacetime metrics $g^{(1)}$ and $g^{(2)}$ and assume that the test particle moves along a geodesic in each branch, we can always find a quantum diffeomorphism that renders the metric flat and thereby definite along the entire path of the test particle. This is because, in each branch of the superposition, there exist Fermi normal coordinates such that the metric is locally flat along the geodesic ([41, Sec. 13.6],[38, 39]). As a result, the metric becomes definite in the region relevant to the computation of the proper time $\tau_{\gamma_{0}}$ (Eq. (2)) and the indefiniteness of causal order is now due to a superposition of worldline crossings. A concrete construction can be found in Appendix A.2. In short, whether ICO is due to delocalized events or a superposition of metrics – a key difference between the optical and gravitational implementation of the quantum switch – depends on the choice of quantum coordinates.
Moreover, note that at least one of the two – the spacetime metric or the location of events – must be in a quantum superposition to verify ICO. There is no way to generate ICO, as defined in Def. 2, in a definite spacetime background with localized events. We can straightforwardly phrase this in terms of a no-go theorem. There is no quantum coordinate system in which the following three statements hold:
@@
This is consistent with the results of [15, Corollary 8.2], despite the different formalisms for ICO.
Operational encoding as quantum mechanical observable — The quantity $s$ characterizing causal order is also an observable in the quantum mechanical sense. To show this, we provide a concrete operational encoding of $s$ in orthogonal qubit states $\mathinner{|{s=\pm 1}\rangle}$ for a setup displaying a superposition of causal orders (see Fig. 3 in the appendix). Assume that the test particle has an internal spin degree of freedom, which precesses around the $z$-axis according to its proper time [42, 43, 44, 45] and is prepared in the initial state $\mathinner{|{b_{0}}\rangle}$ in both branches. In addition, consider the systems $A$ and $B$ to be Alice’s and Bob’s laboratories respectively, with $\gamma_{A}$ and $\gamma_{B}$ their worldlines. We want to ensure that the particle crosses the laboratories at specific proper times. Its first crossing with a laboratory occurs at $\tau^{*}_{1}$, the second crossing at $\tau^{*}_{2}$. Thus, in the first branch, $p_{\mathcal{E}_{A}}^{(1)}=\gamma_{0}(\tau^{*}_{1})$ and $p_{\mathcal{E}_{B}}^{(1)}=\gamma_{0}(\tau^{*}_{2})$, while in the second branch, $p_{\mathcal{E}_{B}}^{(2)}=\gamma_{0}(\tau^{*}_{1})$ and $p_{\mathcal{E}_{A}}^{(2)}=\gamma_{0}(\tau^{*}_{2})$.
When the spin degree of freedom enters the first laboratory at proper time $\tau^{*}_{1}$, it has evolved to be in some state $\mathinner{|{b_{1}}\rangle}$. The crossings of the worldlines are now tuned such that the spin is in the orthogonal state $\mathinner{|{b_{2}}\rangle}$ when it crosses the second laboratory at time $\tau^{*}_{2}$. This ensures that, whenever the particle enters a laboratory, the corresponding agent can measure in the basis $\{\mathinner{|{b_{1}}\rangle},\mathinner{|{b_{2}}\rangle}\}$ without disturbing the state and its time evolution. Upon measurement, each agent encodes the result of their measurement in a memory register with associated Hilbert space $\mathcal{H}_{A,B}\cong\mathbb{C}^{2}$, initialized to $\mathinner{|{0}\rangle}_{A,B}$. Thus, the overall state evolves from the initial state

### v3

# Indefinite causal order and quantum coordinates

@@ Abstract
Classically the causal order of two timelike separated events A and B is fixed – either A before B or B before A. This is no longer true in quantum theory, where it is possible to encounter superpositions of causal orders. The quantum switch is one of the most prominent processes with indefinite causal order. Optical realizations of the quantum switch have been successfully implemented in experiments, but, some argue this merely simulates a process with indefinite causal order and that a superposition of spacetime metrics is required for a true realization. Here, we provide a relativistic definition of causal order between operationally defined events that defines a meaningful observable in both the general relativistic and quantum mechanical sense. Importantly, this observable does not distinguish between the indefinite causal order implemented on an optical bench and the gravitational quantum switch, a gedankenexperiment where the indefinite causal order is achieved by a quantum superposition of gravitational fields. Therefore, our results support the thesis that the optical quantum switch is just as much a realization of indefinite causal order as its gravitational counterpart, which makes use of the quantum mechanical behavior of spacetime.
The study of indefinite causal order (ICO) has seen remarkable progress over the last years both theoretically [1, 2, 3, 4] and experimentally [5, 6, 7, 8, 9, 10]. The field was inspired by the goal to combine general relativity and quantum theory [1] as well as the provable advantage of ICO over causal quantum circuits in quantum information processing [2, 11, 12]. General relativity is a deterministic theory with a dynamical causal structure. Quantum theory is a probabilistic theory with a fixed causal structure. Thus, combining the two would allow for processes to which no definite causal ordering can be assigned.
One of the simplest and most prominent information theoretic primitives of ICO is the quantum switch [2, 3]. Depending on the state of a control qubit, two operations $U_{A}$ and $U_{B}$ are applied to a target system in a superposition of $U_{A}$ before $U_{B}$’ and $U_{B}$ before $U_{A}$’. Concretely, denoting the initial state of the target system by $\mathinner{|{\psi}\rangle}_{T}$ and taking the operations to be unitary, the resulting state is
@@
Various ways of implementing ICO have been reported and described in the literature. In the optical quantum switch, the target system is passing through the setup in a superposition of paths [5, 6, 7]. This type of ICO has been successfully implemented on the optical bench [8, 9, 10]. A different implementation is provided by the gravitational quantum switch [4, 14]. In this gedankenexperiment, a massive object in superposition of two locations sources a gravitational field in superposition, while two agents perform operations on a test particle. A minimal assumption of most quantum gravity theories is that the spacetime metric sourced by a mass in superposition is itself in superposition. This regime, where matter behaves quantum mechanically and spacetime is in a quantum superposition of semi-classical states, has been studied extensively as a possible interface between quantum information theory and quantum gravity (e.g. [15, 16, 17, 18, 19, 4, 20, 21]). Then, choosing an appropriate setup can result in an indefinite causal order due to the different time dilations experienced in each branch of the superposition.
@@
The radically different physics involved in the optical and the gravitational implementations of the quantum switch leads some to question whether it can be claimed that ‘genuine’ or ‘proper’ ICO has been experimentally realized. The quantum information community is roughly divided along two theses [5, 22, 14, 23, 24, 25]. On the one hand, those that claim that the optical switch “can be seen as the first realization of a superposition of causal orders” [5], and on the other, those who view it as a mere simulation of ICO. The latter camp argues that “the current quantum switch experimental implementations do not feature superpositions of causal orders between spacetime events”, and that “these superpositions can only occur in the context of superposed gravitational fields” [14]. This debate is closely intertwined with the question whether realizing ‘genuine’ ICO might be employed in the future to witness non-classicality of spacetime.
@@
To bring clarity into this debate, we cast the notion of ICO in a general relativistic language. This allows us to study the indefiniteness of causal order in a unified setting for both variations of the quantum switch. We first discuss the definition of an event in the context of a superposition of spacetimes and propose an operational, diffeomorphism-invariant notion in terms of worldline coincidences. We then identify a single binary quantity that encodes the causal order of a process, the sign of the proper time, and demonstrate that it is invariant under quantum-controlled diffeomorphisms — a quantum-controlled arbitrary and independent change of coordinates in each branch of the superposition. We further show that this quantity is not only an operationally meaningful observable in the general relativistic but also in the quantum mechanical sense. Crucially, this observable does not differentiate whether ICO is arising due to a superposition of paths on a definite metric or a superposition of spacetimes, that is, it cannot distinguish between the optical and the gravitational quantum switch. This strengthens the claim that the optical and gravitational implementation of the quantum switch exhibit the same ‘type’ of ICO.
@@
This work is motivated by recent research on quantum reference frames [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 20, 38], which suggests that the invariance of physical laws under coordinate transformations extends to quantum superpositions thereof. Indefinite causal order, as we define it here, is a prototype case of a physically meaningful quantity that has no classical counterpart and is an invariant under quantum coordinate changes.
@@
Take two systems and a test particle interacting with each of the systems once. We take each interaction to define an event. The events may correspond to the application of a unitary operation acting on the electron spin. We model this as three timelike curves $\gamma_{0}$, $\gamma_{A}$, and $\gamma_{B}$, with $\gamma_{0}$ coinciding once with $\gamma_{A}$ and once with $\gamma_{B}$ at the coincidences $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$ (see Fig. 1). Coincidences of worldlines define events in a coordinate independent way. Note that while, classically, each such event is associated with a single spacetime point – sometimes referred to as an event in other works – this is will no longer be the case when considering superpositions of spacetimes. Therefore, we reserve the term ‘event’ for events defined in terms of worldline coincidences and refer to spacetime points or manifold points otherwise.
@@
$\tau_{\gamma_{0}}=\frac{1}{c}\int_{p_{\gamma_{0},i}}^{p_{\gamma_{0},f}}\sqrt{-g_{\mu\nu}dx^{\mu}dx^{\nu}},$ (2)
elapsed between two points $p_{\gamma_{0},i}$ and $p_{\gamma_{0},f}$. We take $p_{\gamma_{0},i}$ as the event at which the test particle is released and $p_{\gamma_{0},f}$ as the one at which a final measurement is applied. This fixes a time orientation along $\gamma_{0}$. $\mathcal{E}_{B}$} lies either in the future or past lightcone of $\mathcal{E}_{A}$}. With our time orientation, $\mathcal{E}_{B}$} lies in the future of $\mathcal{E}_{A}$} if $\tau_{A}<\tau_{B}$. This corresponds to being able to signal from $\mathcal{E}_{A}$} to $\mathcal{E}_{B}$}, which is also the operational notion of causal order employed in quantum information. Therefore, the causal order is encoded in the sign of the proper time difference. Define the crossing points $p_{\mathcal{E}_{A}}\equiv\gamma_{0}(\tau_{A})$, $p_{\mathcal{E}_{B}}\equiv\gamma_{0}(\tau_{B})$. The distinction between the events $\mathcal{E}_{X}$ and points $p_{\mathcal{E}_{X}}$ will be important in what follows.
@@
Consider two events $\mathcal{E}_{X}$, $X\in\{A,B\}$, connected by a timelike worldline $\gamma_{0}$ with crossing points $p_{\mathcal{E}_{X}}\equiv\gamma_{0}(\tau_{X})$. The causal order between $\mathcal{E}_{A}$} and $\mathcal{E}_{B}$} is the sign $s\equiv\mathrm{sign}(\Delta\tau)$ of the proper time difference $\Delta\tau\equiv\tau_{B}-\tau_{A}$.
Note that this definition assumes that the events are uniquely defined, i.e. that the worldline $\gamma_{0}$ crosses both $\gamma_{A}$ and $\gamma_{B}$ exactly once.
Events on a superposition of spacetimes — Consider two globally hyperbolic spacetimes $g^{(1)}$ and $g^{(2)}$ each associated to a manifold $\mathcal{M}_{i}$, $i\in\{1,2\}$, diffeomorphic as differentiable manifolds. We take the same setup as above on each spacetime, with two events $\mathcal{E}_{A}^{(i)}$, $\mathcal{E}_{B}^{(i)}$ defined as coincidences. Next, we consider a superposition of two semi-classical, macroscopically distinguishable and thus nearly orthogonal states of the metric, each peaked around $g^{(i)}$ [4, 18, 39]. Formally, using a control qubit whose orthogonal states $\mathinner{|{1}\rangle}$ and $\mathinner{|{2}\rangle}$ distinguish the branches, the quantum state is
@@
with $|\alpha|^{2}+|\beta|^{2}=1$  [^1] Note that Eq. (3) should be understood as suggestive since the Hilbert space assigned to the metric has not properly been introduced. However, if classically distinguishable spacetimes are represented by orthogonal states, one can form linear superpositions thereof and define a suitable inner product in the Hilbert space. They would correspond to the superposition of semi-classical states in a complete theory.. Such a scenario might, for example, arise when placing a gravitating object in a superposition of two locations [15, 16, 17, 18, 19, 4, 20, 21]. Note that the states $\mathinner{|{g^{(i)}}\rangle}$ in Eq. (3) may be thought of as arising from semi-classical states defined at a spacelike surface peaked on initial data for the three–metric and extrinsic curvature such that classically through the equations of motion they generate the four–metrics $g^{(1)}$ and $g^{(2)}$ (cf. [41, Ch. 43]). Moreover, in a covariant formulation, the configurations $g^{(i)}$ can be understood to arise from a stationary phase approximation in the path integral formalism [42].
Next, in order to characterize the causal order between two events on a superposition of spacetimes, we first need to define the notion of an event itself in this context. Since $\mathcal{E}_{X}^{(1)}$ and $\mathcal{E}_{X}^{(2)}$, $X=A,B$ both represent the crossing of the worldline of the test particle with that of system $X$, we posit that they represent the same physical event:
@@
This extends the relativistic notion of coincidences to incorporate superpositions of spacetimes. Note that it differs essentially from the notion of an event as a point on a spacetime manifold. This becomes particularly clear when counting the events involved in our scenario. There are four different manifold-points involved in our scenario – $p_{\mathcal{E}_{A}}^{(1)}$ and $p_{\mathcal{E}_{B}}^{(1)}$ on manifold $\mathcal{M}_{1}$ as well as $p_{\mathcal{E}_{A}}^{(2)}$ and $p_{\mathcal{E}_{B}}^{(2)}$ on manifold $\mathcal{M}_{2}$ (see Fig. 2). However, there are only two distinct physical events $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$ – the crossing of the test particle with system $A$ and the crossing of the test particle with system $B$.
Figure 2: A superposition of two scenarios with different causal orders $s^{(1)}$ and $s^{(2)}$. In each branch, we consider a manifold $\mathcal{M}_{1,2}$, a metric $g^{(1),(2)}$, and the worldlines of three systems. The black, blue, and red line depict $\gamma_{0}$, $\gamma_{A}$, and $\gamma_{B}$, with different transparencies illustrating the superposition. There are two events $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$ defined by the crossing of the test particle with one of the other trajectories. First, we apply a quantum-controlled diffeomorphism $\phi^{(1)}:\mathcal{M}_{1}\to\mathcal{M}$ together with a diffeomorphism $\phi^{(2)}:\mathcal{M}_{2}\to\mathcal{M}$ chosen such that each of the events is associated with a single point on $\mathcal{M}$. The metric and causal order may still be in superposition.
Causal order on a superposition of spacetimes — We can now extend Definition 1 of causal order between events to superpositions of spacetimes.
@@
Thus, the notions of definite and indefinite causal order are conveniently encoded in the product $s^{(1)}s^{(2)}$. Importantly, ICO in the above sense can be due to a superposition of spacetime metrics $g_{\mu\nu}(x)$ or of crossing points $p_{{\mathcal{E}_{A}}}$ and $p_{{\mathcal{E}_{B}}}$ [^3] When considering ICO on a fixed spacetime background due to a superposition of crossing points, we consider two copies of the same manifold and metric, $(\mathcal{M}_{1},g^{(1)})=(\mathcal{M}_{2},g^{(2)})$ in Definition 2.. In the following, we will refer to the latter as ICO due to delocalized events as, in this case, $\mathcal{E}_{A}$} and $\mathcal{E}_{B}$} are assigned different manifold points across the different branches. The above observation follows straightforwardly from Eq. (2), which tells us that the proper time difference depends on both the metric along the curve $\gamma_{0}$ and the coincidences of worldlines that define the events $\mathcal{E}_{A}$} and $\mathcal{E}_{B}$}. Thus, this notion of ICO captures both the optical and the gravitational quantum switch.
@@
This does not mean, of course, that there are no other observables that can distinguish between the two scenarios. Were one to measure the curvature scalar, for example, one would find that it is zero in the optical and non-zero in the gravitational quantum switch in the realization of [4]. Nevertheless, having clarified what we mean by ICO in a general relativistic language, we find that both implementations lead to the same indefiniteness of causal order. In this sense, the optical quantum switch can be seen as just as much a realization of ICO as a hypothetical realization of the gravitational quantum switch. In particular, this implies that quantum spacetime is not a necessary resource for a ‘genuine’ realization of the quantum switch.
@@
Note, however, that whether a process displays ICO ‘due’ to delocalized events or a superposition of spacetime metrics can in general change under quantum diffeomorphisms. To see this, consider a scenario with ICO. Let us assume that we begin in a quantum coordinate system in which ICO is due to delocalized events while the spacetime background is fixed. More specifically, we assume that the crossing points $p_{{\mathcal{E}_{A}}}^{(1)},p_{{\mathcal{E}_{B}}}^{(1)}$ and $p_{{\mathcal{E}_{A}}}^{(2)},p_{{\mathcal{E}_{B}}}^{(2)}$ differ across the branches of the superposition. Now, we can always find a quantum diffeomorphism that maps both configurations onto the same manifold $\mathcal{M}$ while ensuring that the events $\mathcal{E}_{A}$} and $\mathcal{E}_{B}$} are each associated with a single point on this manifold (see also [3]). That is, there always exist diffeomorphisms $\phi^{(1)}:\mathcal{M}_{1}\to\mathcal{M}$ and $\phi^{(2)}:\mathcal{M}_{2}\to\mathcal{M}$ such that
@@
Thus, whether the events $\mathcal{E}_{A}$} and $\mathcal{E}_{B}$} are localized or delocalized depends on the choice of quantum coordinates. If we had defined events in terms of spacetime points (as suggested e.g. in [14]), their number would change under a quantum diffeomorphism. But, as we have seen, the causal order itself is invariant under quantum diffeomorphisms; therefore, the setup must still exhibit ICO. This is because a quantum diffeomorphism also affects the metric. If we start with a definite spacetime metric $g$ and apply a different diffeomorphism in each branch, we end up with a superposition of spacetime metrics $g^{(i)}=(\phi^{(i)})^{-1}_{\ast}(g)$, $i\in\{1,2\}$ [^4] Note that the action of the diffeomorphisms on the fields is the inverse of their actions on the manifold points.. In this sense, a quantum diffeomorphism can shift the indefiniteness due to delocalized events into an indefiniteness of the spacetime metric at localized events. A concrete example of this transformation is provided in Supplemental Material A.1.
Conversely, if we start with a superposition of different spacetime metrics $g^{(1)}$ and $g^{(2)}$ and assume that the test particle moves along a geodesic in each branch, we can always find a quantum diffeomorphism that renders the metric flat and thereby definite along the entire path of the test particle. This is because, in each branch of the superposition, there exist Fermi normal coordinates such that the metric is locally flat along the geodesic ([41, Sec. 13.6],[39, 45]). As a result, the metric becomes definite in the region relevant to the computation of the proper time $\tau_{\gamma_{0}}$ (Eq. (2)) and the indefiniteness of causal order is now due to a superposition of worldline crossings. A concrete construction can be found in Supplemental Material A.2. In short, whether ICO is due to delocalized events or a superposition of metrics – a key difference between the optical and gravitational implementation of the quantum switch – depends on the choice of quantum coordinates. Therefore, in the context of the quantum switch, there is no absolute matter of fact whether ICO is due to quantum features of spacetime.
Moreover, note that at least one of the two – the spacetime metric or the location of events – must be in a quantum superposition to verify ICO. There is no way to generate ICO, as defined in Definition 2, in a definite spacetime background with localized events. We can straightforwardly phrase this in terms of a no-go theorem. There is no quantum coordinate system in which the following three statements hold:
@@
This is consistent with the results of [24, Corollary 1], despite the different formalisms for ICO.
Operational encoding as quantum mechanical observable — The quantity $s$ characterizing causal order is also an observable in the quantum mechanical sense. To show this, we provide a concrete operational encoding of $s$ in orthogonal qubit states $\mathinner{|{s=\pm 1}\rangle}$ for a setup displaying a superposition of causal orders (see Fig. 1 in Supplemental Material B). Assume that the test particle has an internal spin degree of freedom, which precesses around the $z$-axis according to its proper time [47, 48, 49, 50] and is prepared in the initial state $\mathinner{|{b_{0}}\rangle}$ in both branches. In addition, consider the systems $A$ and $B$ to be Alice’s and Bob’s laboratories respectively, with $\gamma_{A}$ and $\gamma_{B}$ their worldlines. We want to ensure that the particle crosses the laboratories at specific proper times. Its first crossing with a laboratory occurs at $\tau^{*}_{1}$, the second crossing at $\tau^{*}_{2}$. Thus, in the first branch, $p_{\mathcal{E}_{A}}^{(1)}=\gamma_{0}(\tau^{*}_{1})$ and $p_{\mathcal{E}_{B}}^{(1)}=\gamma_{0}(\tau^{*}_{2})$, while in the second branch, $p_{\mathcal{E}_{B}}^{(2)}=\gamma_{0}(\tau^{*}_{1})$ and $p_{\mathcal{E}_{A}}^{(2)}=\gamma_{0}(\tau^{*}_{2})$.
When the spin degree of freedom enters the first laboratory at proper time $\tau^{*}_{1}$, it has evolved to be in some state $\mathinner{|{b_{1}}\rangle}$. The crossings of the worldlines are now tuned such that the spin is in the orthogonal state $\mathinner{|{b_{2}}\rangle}$ when it crosses the second laboratory at time $\tau^{*}_{2}$. This ensures that, whenever the particle enters a laboratory, the corresponding agent can measure in the basis $\{\mathinner{|{b_{1}}\rangle},\mathinner{|{b_{2}}\rangle}\}$ without disturbing the state and its time evolution. Upon measurement, each agent encodes either $\tau_{1}^{\ast}$ or $\tau_{2}^{\ast}$, depending on the result of their measurement $b_{1}$ or $b_{2}$ respectively, in a memory register with associated Hilbert space $\mathcal{H}_{A,B}\cong\mathbb{C}^{2}$, initialized to $\mathinner{|{0}\rangle}_{A,B}$. Thus, the overall state evolves from the initial state